# A preorder on the set of links with applications to symmetric unions

### v2

@@ 6. $\pi$-minimal links
If $\frac{1}{\alpha^{\prime}_{1}}+\frac{1}{\alpha^{\prime}_{2}}+\frac{1}{\alpha^{\prime}_{3}}\geq 1$, then $(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})\in\{(2,3,3),(2,3,5),(3,3,3)\}$ since at most one of the $\alpha^{\prime}_{i}$ is even. The triple $\{2,3,3\}$ is not possibles since the order $12$ of $T(2,3,3)$ is not divisible by three distinct prime numbers $\alpha_{1},\alpha_{2},\alpha_{3}$. The triple $\{3,3,3\}$ is not possible also since the torsion elements in the Euclidean triangle group $T(3,3,3)$ have all the same order $3$. If $\{\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3}\}=\{2,3,5\}$, then the only prime divisors of the order $60$ of $T(2,3,5)$ are $2,3$ and $5$. Therefore, $\{\alpha_{1},\alpha_{2},\alpha_{3}\}=\{2,3,5\}$ and $K$ is an elliptic Montesinos knot, which is not the case in (2). In the case (1), $K$ is the $(3,5)$-torus knot and $\Sigma_{2}(K)$ is the Poincaré homology sphere whose fundamental group is the binary icosahedral group $I^{\ast}$ of order $120$. Then $M$ is a Seifert fibered integral homology sphere with finite fundamental group, and hence it is the Poincaré homology sphere and $\pi_{1}(M)$ is isomorphic to $\pi_{1}(\Sigma_{2}(K))$. In particular, the epimorphism $\varphi$ is injective because the groups have the same order.
@@ 7.1. Symmetric unions of knots
Proposition 7.2 can be restated as: if $K_{D}$ is a partial knot for a symmetric union presentation of $K$, then $K\succeq K_{D}$. Then the following result is a corollary of Theorem 1.5. It puts some strong restriction on the partial knot types of the symmetric union presentations for some particular families of knots.
@@
Note that (1) was already proved in [42, Theorem 3.5]. Regarding (2), $\det K=(\det K_{D})^{2}\neq 1$ implies $K_{D}$ is knotted, and hence a $2$-bridge knot, which can be found in [43, Proposition 3.1]. Furthermore, the symmetric union is skew because an epimorphism of a $2$-bridge knot group onto the group of a non-trivial knot never kills the longitude by [4, Propositions 1.8 and 1.10].
@@
For a $2$-bridge symmetric union, the finiteness of the possible partial knots has been recently proved by Lamm and Tanaka [43] by using the fact that the partial knots are alternating because they have at most two bridges.
@@
The epimorphism $\varphi_{D}^{\mathrm{orb}}\colon G^{\mathrm{orb}}(K)\twoheadrightarrow G^{\mathrm{orb}}(K_{D})$ induces an epimorphism $\tilde{\varphi}_{D}\colon\pi_{1}(\Sigma_{2}(K))\twoheadrightarrow\pi_{1}(\Sigma_{2}(K_{D}))$. Therefore, one gets the first inequality in the statement.
@@
Let assume that the knot $11a_{103}$ $\pi$-dominates a knot $K$. Since $11a_{103}$ is a Montesinos knot with three rational tangles, by Corollary 5.4, $K$ is either a knot with $\leq 2$ bridges or a Montesinos knot with three rational tangles.
Suppose that $K$ is a Montesinos knot $K(\frac{\beta_{1}}{\alpha_{1}},\frac{\beta_{2}}{\alpha_{2}},\frac{\beta_{3}}{\alpha_{3}})$. Then the epimorphism $\varphi\colon G^{\mathrm{orb}}(11a_{103})\twoheadrightarrow G^{\mathrm{orb}}(K)$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(11a_{103}))\twoheadrightarrow\pi_{1}(\Sigma_{2}(K))$. Here, the $2$-fold cover $\Sigma_{2}(K)=V(0;e_{0};\frac{\beta_{1}}{\alpha_{1}},\frac{\beta_{2}}{\alpha_{2}},\frac{\beta_{3}}{\alpha_{3}})$ is a Seifert fibered manifold with base $S^{2}(\alpha_{1},\alpha_{2},\alpha_{3})$ and rational Euler number $e_{0}=\sum_{i=1}^{3}\frac{\beta_{i}}{\alpha_{i}}=\frac{|H_{1}(\Sigma_{2}(K);\mathbb{Z})|}{\alpha_{1}\alpha_{2}\alpha_{3}}$. In the same way, $\Sigma_{2}(11a_{103})=V(0;\frac{9}{7};\frac{2}{3},\frac{1}{3},\frac{2}{7})$ is Seifert fibered with base $S^{2}(3,3,7)$ and rational Euler number $\frac{9}{7}=\frac{|H_{1}(\Sigma_{2}(K);\mathbb{Z})|}{63}$, and hence $|H_{1}(\Sigma_{2}(K);\mathbb{Z})|=81$.
@@
Suppose, to the contrary, that there is an epimorphism $\tilde{\varphi}\colon\pi_{1}(M)\twoheadrightarrow\pi_{1}(N)$. Since $\beta_{2}$ is coprime to $3$, $|H_{1}(N;\mathbb{Z})|=21\beta_{1}+14\beta_{2}+6\beta_{3}$ is coprime to $3$. Moreover, $|H_{1}(N;\mathbb{Z})|$ must divide $|H_{1}(M;\mathbb{Z})|=81$, and hence $|H_{1}(N;\mathbb{Z})|=1$. So $N$ is the Brieskorn homology sphere $V(0;\frac{1}{42};\frac{1}{2},\frac{-1}{3},\frac{-1}{7})$, see [35, Theorem 7.12]. Since $M$ and $N$ are aspherical, the epimorphism $\tilde{\varphi}$ can be realized by a map $f\colon M\to N$. Since both $\pi_{1}(M)$ and $\pi_{1}(N)$ have rank $2$, the $\pi_{1}$-surjective map $f$ is of non-zero degree by [56, Theorem 2.1]. Then, by [58, Theorem 3.2 and Remark 2], $f$ is homotopic to $g\circ\pi$, where $\pi$ is a composition of finitely many vertical pinches and $g$ is a fiber-preserving branched covering. Since $M$ has only three exceptional fibers, a vertical pinch on $M$ would produce a Seifert fibered manifold with $\leq 2$ exceptional fibers, and hence with cyclic fundamental group. Therefore, up to homotopy, $f$ cannot factorize through a vertical pinch since $\pi_{1}(N)$ is not cyclic. It follows that $f$ is homotopic to a fiber-preserving branched covering $g$. Let $d_{1}$ be the degree of $g$ along the fiber and $d_{2}$ the degree of the induced map $\bar{g}\colon S^{2}(3,3,7)\to S^{2}(2,3,3)$. Then, $\frac{4}{21}=|\chi(S^{2}(3,3,7))|\geq d_{2}|\chi(S^{2}(2,3,7))|=\frac{d_{2}}{42}$. This implies that $8\geq d_{2}$. Moreover, $\frac{9}{7}=e_{0}(M)=\frac{d_{2}}{d_{1}}e_{0}(N)=\frac{d_{2}}{42d_{1}}$ by [52, Theorem 1.2]. Then $d_{1}=\frac{d_{2}}{54}\leq\frac{8}{54}$, which is impossible. ∎
@@
The proofs follow from Proposition 7.10 and the following well-known facts. First, $b(K_{D})$ verifies the inequality $b(K_{D})\geq\operatorname{rank}\pi_{1}(\Sigma_{2}(K_{D}))+1$, see [11, Lemma 1.6]. Next, for a Seifert matrix $S$ of $K_{D}$, the sum $S+S^{T}$ gives a presentation matrix for the $\mathbb{Z}$-module $H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$, and therefore $2g(K_{D})\geq\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$. Finally, Wendt’s inequality [63] shows that $u(K_{D})\geq\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$ (see also [38]). ∎
@@ 8. The bridge number and preorder $\succeq$
Let us assume that $K_{1}$ is a $(p_{1},q_{1})$-torus knot and $K_{2}$ a $(p_{2},q_{2})$-torus knot with $3\leq p_{1}<q_{1}$, $3\leq p_{2}<q_{2}$, $p_{1},q_{1}$ coprime, and $p_{2},q_{2}$ coprime. Furthermore, if $p_{i}=3$, then $q_{i}\geq 7$. Since $K_{1}\succeq K_{2}$ there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(K_{1})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$ which, by the proof of Theorem 1.5(4), sends the infinite cyclic center $Z_{1}$ of $G^{\mathrm{orb}}(K_{1})$ into the infinite cyclic center $Z_{2}$ of $G^{\mathrm{orb}}(K_{2})$. The quotients $G(K_{1})/Z_{1}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{1},q_{1}))=T(2,p_{1},q_{1})$ and $G(K_{2})/Z_{2}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{2},q_{2}))=T(2,p_{2},q_{2})$ are hyperbolic triangle groups since $\frac{1}{2}+\frac{1}{p_{i}}+\frac{1}{q_{i}}<1$ for $i=1,2$. In particular, they are centerless, and thus the epimorphism $\varphi$ induces an epimorphism $\bar{\varphi}\colon T(2,p_{1},q_{1})\twoheadrightarrow T(2,p_{2},q_{2})$. Like in the proof of Proposition 6.6, it follows from the presentation of the triangle group $T(2,p_{1},q_{1})$ that each image $\bar{\varphi}(x)$, $\bar{\varphi}(y)$ and $\bar{\varphi}(z)$ is not trivial, otherwise the image $\bar{\varphi}(T(2,p_{1},q_{1}))$ would be trivial. Therefore, the non-trivial elements $a=\bar{\varphi}(x)$ and $b=\bar{\varphi}(y)$ verifies $a^{2}=1$ and $b^{p^{\prime}_{1}}=1$ with $p^{\prime}_{1}$ dividing $p_{1}$. The elements $a$ and $b$ generate the image $\bar{\phi}(T(2,p_{1},q_{1}))=T(2,p_{2},q_{2})$ which is a discrete subgroup of $\mathit{PSL}(2,\mathbb{R})$ since $T(2,p_{2},q_{2})$ is a hyperbolic triangle group. As in the proof of Proposition 6.6, up to taking suitable powers $u=a^{\pm 1}$ and $v=b^{k}$, $k$ coprime to $p^{\prime}_{1}$ to normalize the matrix representatives of $u$ and $v$ in $SL(2,\mathbb{R})$, it follows that at least one of Cases (I)–(VII) in [41, Theorem 2.3] holds. Since the triangle group $T(2,p_{2},q_{2})$ is co-compact, Case (II) is impossible by [41, Figures 2 and 3]. Case (III) cannot happens, otherwise $0=\operatorname{tr}(u)=\operatorname{tr}(v)$ and $\bar{\varphi}(T(2,p_{1},q_{1}))=T(2,\ell,n)$ with $\operatorname{tr}(u)=\operatorname{tr}(v)=\cos(\pi-\frac{\pi}{\ell})$ by [41, Section 3]. Then $\ell=2$, which contradicts $\frac{1}{2}+\frac{1}{\ell}+\frac{1}{n}<1$ because $T(2,\ell,n)=T(2,p_{2},q_{2})$ is a hyperbolic triangle group. If Case (IV) happens, then $u$ and $v$ generate a hyperbolic triangle group $T(2,3,n)$ by [41, Section 3]. Since $T(2,3,n)=T(2,p_{2},q_{2})$ with $3\leq p_{2}<q_{2}$ and $p_{2}$ coprime to $q_{2}$, it follows that $p_{2}=3$ and $q_{2}=n$. Therefore,

### v3

@@ 6. $\pi$-minimal links
If $\frac{1}{\alpha^{\prime}_{1}}+\frac{1}{\alpha^{\prime}_{2}}+\frac{1}{\alpha^{\prime}_{3}}\geq 1$, then $(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})\in\{(2,3,3),(2,3,5),(3,3,3)\}$ since at most one of the $\alpha^{\prime}_{i}$ is even. The triple $\{2,3,3\}$ is not possible since the order $12$ of $T(2,3,3)$ is not divisible by three distinct prime numbers $\alpha_{1},\alpha_{2},\alpha_{3}$. The triple $\{3,3,3\}$ is also not possible since the torsion elements in the Euclidean triangle group $T(3,3,3)$ have all the same order $3$. If $\{\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3}\}=\{2,3,5\}$, then the only prime divisors of the order $60$ of $T(2,3,5)$ are $2,3$ and $5$. Therefore, $\{\alpha_{1},\alpha_{2},\alpha_{3}\}=\{2,3,5\}$ and $K$ is an elliptic Montesinos knot, which is not the case in (2). In the case (1), $K$ is the $(3,5)$-torus knot and $\Sigma_{2}(K)$ is the Poincaré homology sphere whose fundamental group is the binary icosahedral group $I^{\ast}$ of order $120$. Then $M$ is a Seifert fibered integral homology sphere with finite fundamental group, and hence it is the Poincaré homology sphere and $\pi_{1}(M)$ is isomorphic to $\pi_{1}(\Sigma_{2}(K))$. In particular, the epimorphism $\varphi$ is injective because the groups have the same order.
@@ 7.1. Symmetric unions of knots
Using Lemma 2.11, Proposition 7.2 can be restated as: if $K_{D}$ is a partial knot for a symmetric union presentation of $K$, then $K\succeq K_{D}$. Then the following result is a corollary of Theorem 1.4. It puts some strong restriction on the partial knot types of the symmetric union presentations for some particular families of knots.
@@
When $K$ is a $2$-bridge knot, $\det K=(\det K_{D})^{2}\neq 1$, then $K_{D}$ is knotted and hence a $2$-bridge knot. The symmetric union is skew because an epimorphism of a $2$-bridge knot group onto the group of a non-trivial knot never kills the longitude by [4, Propositions 1.8 and 1.10].
@@
For a $2$-bridge symmetric union, the finiteness of the possible partial knots has been recently proved by Lamm and Tanaka [43] by using the fact that the partial knots are alternating because they have at most two bridges. See also [61].
@@
By Lemma 2.10, the epimorphism $\varphi_{D}^{\mathrm{orb}}\colon G^{\mathrm{orb}}(K)\twoheadrightarrow G^{\mathrm{orb}}(K_{D})$ induces an epimorphism $\tilde{\varphi}_{D}\colon\pi_{1}(\Sigma_{2}(K))\twoheadrightarrow\pi_{1}(\Sigma_{2}(K_{D}))$. Therefore, one gets the first inequality in the statement.
@@
Let us assume that the knot $11a_{103}$ $\pi$-dominates a knot $K$. Since $11a_{103}$ is a Montesinos knot with three rational tangles, by Corollary 5.4, $K$ is either a knot with $\leq 2$ bridges or a Montesinos knot with three rational tangles.
Suppose that $K$ is a Montesinos knot $K(\frac{\beta_{1}}{\alpha_{1}},\frac{\beta_{2}}{\alpha_{2}},\frac{\beta_{3}}{\alpha_{3}})$. Then, by Lemma 2.10, the epimorphism $\varphi\colon G^{\mathrm{orb}}(11a_{103})\twoheadrightarrow G^{\mathrm{orb}}(K)$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(11a_{103}))\twoheadrightarrow\pi_{1}(\Sigma_{2}(K))$. Here, the $2$-fold cover $\Sigma_{2}(K)=V(0;e_{0};\frac{\beta_{1}}{\alpha_{1}},\frac{\beta_{2}}{\alpha_{2}},\frac{\beta_{3}}{\alpha_{3}})$ is a Seifert fibered manifold with base $S^{2}(\alpha_{1},\alpha_{2},\alpha_{3})$ and rational Euler number $e_{0}=\sum_{i=1}^{3}\frac{\beta_{i}}{\alpha_{i}}=\frac{|H_{1}(\Sigma_{2}(K);\mathbb{Z})|}{\alpha_{1}\alpha_{2}\alpha_{3}}$. In the same way, $\Sigma_{2}(11a_{103})=V(0;\frac{9}{7};\frac{2}{3},\frac{1}{3},\frac{2}{7})$ is Seifert fibered with base $S^{2}(3,3,7)$ and rational Euler number $\frac{9}{7}=\frac{|H_{1}(\Sigma_{2}(K);\mathbb{Z})|}{63}$, and hence $|H_{1}(\Sigma_{2}(K);\mathbb{Z})|=81$.
@@
Suppose, to the contrary, that there is an epimorphism $\tilde{\varphi}\colon\pi_{1}(M)\twoheadrightarrow\pi_{1}(N)$. Since $\beta_{2}$ is coprime to $3$, $|H_{1}(N;\mathbb{Z})|=21\beta_{1}+14\beta_{2}+6\beta_{3}$ is coprime to $3$. Moreover, $|H_{1}(N;\mathbb{Z})|$ must divide $|H_{1}(M;\mathbb{Z})|=81$, and hence $|H_{1}(N;\mathbb{Z})|=1$. So $N$ is the Brieskorn homology sphere $V(0;\frac{1}{42};\frac{1}{2},\frac{-1}{3},\frac{-1}{7})$, see [35, Theorem 7.12]. Since $M$ and $N$ are aspherical, the epimorphism $\tilde{\varphi}$ can be realized by a map $f\colon M\to N$. Since both $\pi_{1}(M)$ and $\pi_{1}(N)$ have rank $2$, the $\pi_{1}$-surjective map $f$ is of non-zero degree by [56, Theorem 2.1]. Then, by [58, Theorem 3.2 and Remark 2], $f$ is homotopic to $g\circ\pi$, where $\pi$ is a composition of finitely many vertical pinches and $g$ is a fiber-preserving branched covering. Since $M$ has only three exceptional fibers, a vertical pinch on $M$ would produce a Seifert fibered manifold with $\leq 2$ exceptional fibers, and hence with cyclic fundamental group. Therefore, up to homotopy, $f$ cannot factorize through a vertical pinch since $\pi_{1}(N)$ is not cyclic. It follows that $f$ is homotopic to a fiber-preserving branched covering $g$. Let $d_{1}$ be the degree of $g$ along the fiber and $d_{2}$ the degree of the induced map $\bar{g}\colon S^{2}(3,3,7)\to S^{2}(2,3,7)$. Then, $\frac{4}{21}=|\chi(S^{2}(3,3,7))|\geq d_{2}|\chi(S^{2}(2,3,7))|=\frac{d_{2}}{42}$. This implies that $8\geq d_{2}$. Moreover, $\frac{9}{7}=e_{0}(M)=\frac{d_{2}}{d_{1}}e_{0}(N)=\frac{d_{2}}{42d_{1}}$ by [52, Theorem 1.2]. Then $d_{1}=\frac{d_{2}}{54}\leq\frac{8}{54}$, which is impossible. ∎
@@
The proofs follow from Proposition 7.10 and the following well-known facts. First, $b(K_{D})$ satisfies the inequality $b(K_{D})\geq\operatorname{rank}\pi_{1}(\Sigma_{2}(K_{D}))+1$, see [11, Lemma 1.6]. Next, for a Seifert matrix $S$ of $K_{D}$, the sum $S+S^{T}$ gives a presentation matrix for the $\mathbb{Z}$-module $H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$, and therefore $2g(K_{D})\geq\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$. Finally, Wendt’s inequality [64] shows that $u(K_{D})\geq\operatorname{rank}H_{1}(\Sigma_{2}(K_{D});\mathbb{Z})$ (see also [38]). ∎
@@ 8. The bridge number and preorder $\succeq$
Let us assume that $K_{1}$ is a $(p_{1},q_{1})$-torus knot and $K_{2}$ a $(p_{2},q_{2})$-torus knot with $3\leq p_{1}<q_{1}$, $3\leq p_{2}<q_{2}$, $p_{1},q_{1}$ coprime, and $p_{2},q_{2}$ coprime. Furthermore, if $p_{i}=3$, then $q_{i}\geq 7$. Since $K_{1}\succeq K_{2}$ there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(K_{1})\twoheadrightarrow G^{\mathrm{orb}}(K_{2})$ which, by the proof of Theorem 1.4(4), sends the infinite cyclic center $Z_{1}$ of $G^{\mathrm{orb}}(K_{1})$ into the infinite cyclic center $Z_{2}$ of $G^{\mathrm{orb}}(K_{2})$. The quotients $G(K_{1})/Z_{1}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{1},q_{1}))=T(2,p_{1},q_{1})$ and $G(K_{2})/Z_{2}\cong\pi_{1}^{\mathrm{orb}}(S^{2}(2,p_{2},q_{2}))=T(2,p_{2},q_{2})$ are hyperbolic triangle groups since $\frac{1}{2}+\frac{1}{p_{i}}+\frac{1}{q_{i}}<1$ for $i=1,2$. In particular, they are centerless, and thus the epimorphism $\varphi$ induces an epimorphism $\bar{\varphi}\colon T(2,p_{1},q_{1})\twoheadrightarrow T(2,p_{2},q_{2})$. Like in the proof of Proposition 6.6, it follows from the presentation of the triangle group $T(2,p_{1},q_{1})$ that each image $\bar{\varphi}(x)$, $\bar{\varphi}(y)$ and $\bar{\varphi}(z)$ is not trivial, otherwise the image $\bar{\varphi}(T(2,p_{1},q_{1}))$ would be trivial. Therefore, the non-trivial elements $a=\bar{\varphi}(x)$ and $b=\bar{\varphi}(y)$ satisfies $a^{2}=1$ and $b^{p^{\prime}_{1}}=1$ with $p^{\prime}_{1}$ dividing $p_{1}$. The elements $a$ and $b$ generate the image $\bar{\phi}(T(2,p_{1},q_{1}))=T(2,p_{2},q_{2})$ which is a discrete subgroup of $\mathit{PSL}(2,\mathbb{R})$ since $T(2,p_{2},q_{2})$ is a hyperbolic triangle group. As in the proof of Proposition 6.6, up to taking suitable powers $u=a^{\pm 1}$ and $v=b^{k}$, $k$ coprime to $p^{\prime}_{1}$ to normalize the matrix representatives of $u$ and $v$ in $SL(2,\mathbb{R})$, it follows that at least one of Cases (I)–(VII) in [41, Theorem 2.3] holds. Since the triangle group $T(2,p_{2},q_{2})$ is co-compact, Case (II) is impossible by [41, Figures 2 and 3]. Case (III) cannot happens, otherwise $0=\operatorname{tr}(u)=\operatorname{tr}(v)$ and $\bar{\varphi}(T(2,p_{1},q_{1}))=T(2,\ell,n)$ with $\operatorname{tr}(u)=\operatorname{tr}(v)=\cos(\pi-\frac{\pi}{\ell})$ by [41, Section 3]. Then $\ell=2$, which contradicts $\frac{1}{2}+\frac{1}{\ell}+\frac{1}{n}<1$ because $T(2,\ell,n)=T(2,p_{2},q_{2})$ is a hyperbolic triangle group. If Case (IV) happens, then $u$ and $v$ generate a hyperbolic triangle group $T(2,3,n)$ by [41, Section 3]. Since $T(2,3,n)=T(2,p_{2},q_{2})$ with $3\leq p_{2}<q_{2}$ and $p_{2}$ coprime to $q_{2}$, it follows that $p_{2}=3$ and $q_{2}=n$. Therefore,